An investor has a risk aversion coefficient of 5. The expected return and standard deviation of...

  1. An investor has a risk aversion coefficient of 5. The expected return and standard deviation of the optimal risky portfolio are 15% and 25%, respectively. If the Sharpe ratio of the optimal capital allocation line is 0.48, what is the proportion of the investor’s combined portfolio that should be invested in the risky portfolio that would maximise their utility?
Answer #1

U = E( r ) - 0.5A012 y* = (E(rp) - rf)/Aop^2 Here E(rp) 15% 25% op Sharpe ratio (Elrp) - rf)/op 0.48 = (.15-rf)/.25 rf = 3.00